XYZ LTD is considering replacing a machine with the following cash flows and salvage values
year cash flow(000) Salvage value (000)
1 5,000 6,000
2 3,800 3,000
3 2,500 1,000
the company project requires an initial cash investment of 7.5 million and cost of capital is 12 %
advice management when to replace asset with
a) identical asset
b) Non- identical asset
1
Expert's answer
2019-08-05T12:01:06-0400

year 1= 6000- 5000=1000

year 2= 2 3000-3800= -800

year 3= 1000- 2500= -1500

Decision Rule using NPV

The decision rule under NPV is to:

Accept the project if the NPV is positive

 Reject the project if NPV is negative

If this value is positive, the project is profitable and viable. If this value is negative, the project is loss-making and should be avoided.

In this case the NPV is negative so the is loss making and the company should

a. The asset should be replaced with idetical asset in the second year sice the NPV is positive, 1000-800=200

b. The asset should be replaced with non-identical asset in the third year sice the NPV is negative, 1000-800-1500=-1700